OpenStudy (airyana1114):

Let f(x)=x-2 and g(x)=x^2-7x-9. Find f(g(-1)). Can someone help me step by step?

OpenStudy (jdoe0001):

\(\bf g(x)=x^2-7x-9\qquad \qquad g(-1)=?\)

OpenStudy (airyana1114):

Is it -15? @jdoe0001

OpenStudy (anonymous):

Yes, it is -15 :)

OpenStudy (airyana1114):

Do I just plug -15 in for x in f(x)? @Tommynaiter

OpenStudy (anonymous):

Yes, you are correct.

OpenStudy (airyana1114):

Awesome, thank you guys :) @Tommynaiter @jdoe0001

OpenStudy (jdoe0001):

hmmmm what's hmmm \(\Large (-1)^2=?\)

OpenStudy (airyana1114):

1 @jdoe0001

OpenStudy (jdoe0001):

\(\bf g(x)=x^2-7x-9\qquad \qquad g(-1)=(-1)^2-7(-1)-9 \\ \quad \\ g(-1)=1+7-9\implies g(-1)=8-9\implies g(-1)=-1\)

OpenStudy (jdoe0001):

and yes, then you just plug in g(-1) into f(x)

OpenStudy (airyana1114):

Oh okay, I accidentally did 7(-1) not -7(-1). I mixed up my signs. Okay, thank you :) So my final answer would be -3 correct? @jdoe0001